Hockey teams receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 63 points. They won 9 more games than they tied. how many wins and how many ties did the team have?
Found 2 solutions by tinkerbell146, richwmiller:
Answer by tinkerbell146(2)

w=wins-2 point
t=ties-1 point
w+t=63
t=9+w
Hope this helps

Answer by richwmiller(17219)

2w+t=63
w=t+9
the new tutor forgot that there are 2 points for a win and 1 for a tie
t=15 w=24
check
24=15+9
ok
2*24+15=63
48+15=63
ok